Circular No. 8693

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006ba, 2006bb, 2006bc, 2006bd, 2006be, 2006bf Six new apparent supernovae have been reported from unfiltered CCD

images: 2006ba by L. A. G. Monard (cf. IAUC 8690); 2006bc by R. Martin (cf. IAUC 8496); 2006be by N. J. Ponticello and W. Li (LOSS/KAIT; cf. IAUC 8691); and the rest by T. Puckett (cf. IAUC 8691; with R. Gagliano for 2006bb and 2006bf, and with L. Cox for 2006bd).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006ba Mar. 19.81 9^{h}43^{m}13.^{s}40 − 9^{o}36′53′′.0 18.4 21′′ E, 8′′ S
2006bb Mar. 25.10 8 33 31.09 +41 31 04.1 17.4 3′′.6 E, 27′′.5 S
2006bc Mar. 24.65 7 21 16.50 −68 59 57.3 16.0 19′′ W, 8′′ N
2006bd Mar. 26.20 11 38 28.46 +20 31 34.4 18.2 14′′.5 W, 5′′.2 S
2006be Mar. 24.51 15 45 39.00 +28 05 19.2 15.4 6′′.2 W, 0′′.5 S
2006bf Mar. 27.35 12 58 50.68 + 9 39 30.1 17.7 3′′.1 W, 15′′.5 N

Additional approximate magnitudes for 2006ba in NGC 2980: Feb. 5.036 UT, [18.8; Mar. 25.847, 17.7 ± 0.2. Nothing is visible at the location of 2006ba on a Digitized Sky Survey image (limiting red mag 20.5). Additional magnitudes for 2006bb in UGC 4468: Mar. 7, [18.5; 26.021, 17.3. Additional approximate magnitudes for 2006bc in NGC 2397: Mar. 16.587, [18.5; 25.618, 16.0. Additional magnitudes for 2006bd in UGC 6609: Jan. 5, [19.2; Mar. 27.208, 18.0. Additional KAIT magnitudes for 2006be in IC 4582: Feb. 22.53, [19.0; Mar. 27.45, 15.4. Additional magnitudes for 2006bf in UGC 8093: Jan. 6, [19.3; Mar. 28.117, 17.7.

COMET 73P/SCHWASSMANN-WACHMANN Further to CBET 439 and IAUC 8692, E. J. Christensen, Mt. Lemmon

Survey, has reported four further components of comet 73P from observations with the 1.5-m reflector on Mar. 24.5 and 25.5 UT (and in some cases from Mar. 23.5). Component ‘P’, with magnitude estimates in the range 19.8–21.7, has positions corresponding to T = 2006 June 7.88 TT. Component ‘Q’, of mag 20.7–21.3, has T = 2006 June 7.74. Component ‘R’, of mag 20.5–20.8, has T = 2006 June 8.20. Component ‘S’, of mag 20.9–21.5, has T only some 0.001 day earlier than component ‘K’, which has now been redetected at mag 20.9–21.3. The astrometric observations are provided on MPEC 2006-F33.

Visual total-magnitude estimates for component ‘C’: Mar. 19.96 UT, 10.7 (W. Hasubick, Buchloe, Germany, 0.44-m reflector); 23.80, 10.5 (Y. Nagai, Gunma, Japan, 45×141 binoculars).

2006 March 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green